OpenStudy (anonymous):

use the trigonometric identities to answer the following : 5. 1-sin^2x= 6. sec^2x-1=

OpenStudy (loser66):

or google

OpenStudy (anonymous):

here's a fun trick: remember only this one. for life. \[\sin^2 (x) + \cos^2(x) = 1\] and if you divide everything by cos^2(x) \[\frac{ \sin^2(x) }{ \cos^2(x) } + \frac{ \cos^2(x) }{ \cos^2(x) } = \frac{ 1 }{ \cos^2(x) }\]\[\tan^2(x) + 1 = \sec^2(x)\]

OpenStudy (jhannybean):

So now you can play around with these identities given to you by @jdoe0001 and @Euler271 to solve for the function you're looking for.